# Stabilizing the Splits through Minimax Decision Trees

### v2

@@ D.1 Performance metrics
D.2 Higher-dimensional input domain
@@
Sample Size Dimension Sklearn VarianceSplit MinimaxSplit
@@
Table 1 compares three decision tree methods—scikit-learn’s standard implementation, VarianceSplit and MinimaxSplit—across various sample sizes $n$ and dimensions $d$ for the Powell function. In the dense design regime ($n>d$), we observe relatively consistent performance across all three methods. However, as we transition into sparse designs ($d>n$), more pronounced differences emerge. The MinimaxSplit method often demonstrates superior performance in these high-dimensional, data-scarce scenarios. For instance, with 10 samples and 256 dimensions, the MinimaxSplit approach achieves an MSE of 8.91 × 104, outperforming both scikit-learn (1.07 × 105) and the VarianceSplit method (9.58 × 104).
Appendix E Pseudocode for algorithms
@@
Figure 12: Comparison of node-level split selection in two random forest variants. Left: a random-dimension random forest selects a random feature uniformly for each node and choose the best split over the chosen feature. Right: a regular RF node samples $m_{\mathrm{try}}$ coordinates uniformly from the $d$ features and chooses the best split over all features according to the chosen impurity/criterion (the highlighted feature illustrates the winner).

### v3

@@ E.1 Performance metrics
E.2 Two-dimensional input domain
For this experiment, we use a synthetic dataset generated from a complex, non-linear function of two features, $\mathbf{x}=(x_{1},x_{2})$. The true function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ used for generating the target values is defined as:
$\begin{split}f(\mathbf{x})&=\left(2-2.1u_{1}^{2}+\frac{u_{1}^{4}}{3}\right)u_{1}^{2}+u_{1}u_{2}+\left(-4+4u_{2}^{2}\right)u_{2}^{2};\\ &u_{1}=3(x_{1}-0.5),\\ &u_{2}=3(x_{2}-0.5),\end{split}$ (76)
where $x_{1},x_{2}$ are independently and uniformly sampled from $[0,1]$. This function incorporates both quadratic and higher-order polynomial terms, creating a complex surface with multiple peaks and valleys. The visualizations of the prediction surfaces, as shown in Figures 15 and 16, further illustrate the improved accuracy and smoother transitions in the predicted values when using the proposed methods. In addition to decision trees fitted with MSE ($L^{2}$) loss with these three kinds of splitting criteria, we also provide trees fitted with $L^{1}$ loss for completeness.
Figure 15: Predictions and true function values of (76), visualized using heatmaps. Top Row: The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models. The next three subplots depict the predictions of the variance-based decision tree, minimax decision tree, and cyclic minimax decision tree (depth $k=10$), respectively. Each plot is annotated with the corresponding MSE to quantitatively assess the accuracy. Bottom Row: The final four subplots visualize the predictions from four different random forest models (using Algorithm 1 of Appendix F), each built with varying combinations of the three error methods. Again, each plot is labeled with the MSE to facilitate direct comparison.
Figure 16: The predictions and true function values are visualized using heatmaps, allowing for a clear comparison of how well each model approximates the underlying function (76). The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models. The next three subplots depict the predictions of the $L^{1}$ VarianceSplit, $L^{1}$ MinimaxSplit, $L^{2}$ VarianceSplit, and $L^{2}$ MinimaxSplit. Top Row: maximum depth 2; Middle Row: maximum depth 6; Bottom Row: maximum depth 10.
In this example (as well as in Sections 5.3 and 3 below), we also incorporate an $L^{1}$ variation of the VarianceSplit and MinimaxSplit algorithms, where in the splitting rule, we replace the variances in (1), (4), and (5) by the $L^{1}$ distance from the mean. For instance, the $L^{1}$ variation of (1) is
$\begin{split}(j,x_{j})&=\operatorname*{arg\,min}_{(j,\,x_{j})}\Big(\mathbb{E}\big[|Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}]|{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}\big]\\ &\hskip 56.9055pt+\mathbb{E}\big[|Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}]|{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}\big]\Big).\end{split}$ (77)
The optimal $L^{1}$ split point may not be unique, in which case we break ties arbitrarily.
E.3 Higher-dimensional input domain
@@
Sample Size Dimension Scikit-learn VarianceSplit MinimaxSplit
@@
Table 1 compares three decision tree methods—scikit-learn’s standard implementation, VarianceSplit and MinimaxSplit—across various sample sizes $n$ and dimensions $d$ for the Powell function. In the dense design regime ($n>d$), we observe relatively consistent performance across all three methods. However, as we transition into sparse designs ($d>n$), more pronounced differences emerge. The MinimaxSplit method often demonstrates superior performance in these high-dimensional, data-scarce scenarios. For instance, with 10 samples and 256 dimensions, the MinimaxSplit approach achieves an MSE of 1.14 × 105, outperforming both scikit-learn (1.24 × 105) and the VarianceSplit method (1.35 × 105).
Appendix F Pseudocode for algorithms
@@
Figure 17: Comparison of node-level split selection in two random forest variants. Left: a random-dimension random forest selects a random feature uniformly for each node and chooses the best split over the chosen feature. Right: a regular RF node samples $m_{\mathrm{try}}$ coordinates uniformly from the $d$ features and chooses the best split over all sampled features according to the chosen impurity/criterion (the highlighted feature illustrates the winner).